• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Mauro Iodice

    Log entry time 08:06:32 on June10,2000

    Entry number 43349

    keyword=owl shift summary - June 10

    00:10 Q3 Electron arm quenched. It is now reset (from downstairs).
    00:20 Start good run #1869 - Always kinematic 155,180
    01:40 Q3 Electron arm QUENCHED AGAIN - END Run 1873
    Paged Stevens (tech. on call) - He will come here. Now we have asked for
    an access to reset again Q3 from the Hall.
    02:15 Q3 reset. Cycling ... setting
    02:28 START taking data : run 1874
    02:30 The RF feedback was off (only for two minutes in this run). Asked to put it on.
    03:07 After some minutes of beam off, MCC told us they are having problems with a RF cavity. Run 1876 ended .
    03:27 Beam is back. Start run 1877
    03:50 The beam is not stable. In average, every 2 minutes there is a trip (lasting
    about 1min)
    04:03 Run 1878 ended because of RF cavity problems
    04:12 Beam is back
    05:00 MCC called. They want to work on the cavities. The beam will be down for about ten minutes
    05:20 Q3 QUENCHED AGAIN. Mark Stevens actually told us (at around 2:00) he would have came here, but he didn't. We asked an access to reset the quad
    05:35 Q3 reset, coming up.
    05:51 START Run 1882
    07:10 As a rough estimate we have collected about 440K true coincidences during this shift (runnr 1869:1886)
    08:00 Mark Stevens came to fix the Q3 problem.


    StartTime.......Run#...........#events.............comment
    00:19...............1869............1.3M.................good
    00:32...............1870............1.3M.................good
    00:48...............1871............1.3M.................good
    01:02...............1872............1.3M.................good
    01:18...............1873............1.2M.................Ended for Q3 QUENCH
    02:28...............1874............1.3M.................good
    02:41...............1875............1.3M.................good
    02:54...............1876............0.3M.................good Ended for RF Cavity problems
    03:27...............1877............1.3M.................good
    03:43...............1878............0.6M.................good Ended for RF Cavity problems
    04:13...............1879............1.3M.................good
    04:28...............1880............1.3M.................good
    04:44...............1881............0.7M.................good Ended to allow work on cavities
    05:51...............1882............1.3M.................good
    06:09...............1883............1.3M.................good
    06:24...............1884............1.3M.................good
    06:39...............1885............1.3M.................good
    06:53...............1886............1.3M.................good
    07:10...............1887............1.3M.................good
    07:25...............1888............1.3M.................good
    07:47...............1889............1.3M.................good
    08:02...............1890 .....running. Left to the next shift!